OpenStudy (anonymous):

******WILL FAN AND MEDAL!****** Which set of side lengths will form a right triangle? 5, 11, 13 9, 24, 25 7, 24, 25 8, 13, 17

OpenStudy (idku):

has to be a^2+b^2=c^2

OpenStudy (anonymous):

What do you mean by that?

OpenStudy (idku):

plug in the biggest side for c, and smaller sides for a and b

OpenStudy (idku):

triangle with sides a b and c is only a right triangle if: a^2+b^2=c^2

OpenStudy (anonymous):

So I put in the answer choices with power to the 2nd?

OpenStudy (anonymous):

and add them up? @idku

OpenStudy (anonymous):

@LynFran @Falling_In_Katt

OpenStudy (anonymous):

@aleeah1404 @Ddaddysgirl14 @math1234 @mathmate @Rabakku @rvc @TillLindemann @welshfella

OpenStudy (anonymous):

I say a or b cause the bottom is the smallestr and the 2 sides are the same

OpenStudy (anonymous):

Okay.

OpenStudy (welshfella):

try the first one the longest side = 13 13^2 = 169 12^2 = 144 5^2 = 25 what do you infer from these results?

OpenStudy (anonymous):

What do you mean by infer?

OpenStudy (welshfella):

what conclusion do you come to? are these the 3 sides of a right angled triangle or not ?

OpenStudy (anonymous):

d

OpenStudy (anonymous):

Oh, I'm guessing it is.... I mean I kind of see what you did, and I can tell that the 3 lengths equaled to the right angle... I guess..

OpenStudy (welshfella):

well 13^2 = 12^2 + 5^2 - which can only mean that the triangle is right angled - according to the Pythagoras theorem

OpenStudy (welshfella):

the square on the hypotenuse ( largest side) = sum of the squares on the other 2 sides - is only true for right angled triangles

OpenStudy (anonymous):

Okay.

OpenStudy (welshfella):

in symbols that is c^2 = a^2 + b^2 where c is the hypotenuse

OpenStudy (anonymous):

Okay. Thank you!

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

you might like to check out the other 3 - theres at least one which is a right angled triangle

OpenStudy (anonymous):

Okay.